Opened 9 years ago

Closed 7 years ago

Last modified 3 months ago

#371 closed flight processing (fixed)

EUFAR10/01 (AIMWETLAB), flight day 238b/2010, Lake Balaton Peninsula

Reported by: iopa Owned by: knpa
Priority: alpha 4 high Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by iopa)

Data location: ~arsf/workspace/EU10_03-2010_238b_Lake_Balaton_Peninsula

Data arrived from ARSF via SATA disk Q on 15/09/2010.

Scientific objective: Mapping of species composition and stand health of shore vegetation, mapping of shore elevation and inundation structure.

This flight will also be used for EUFAR10-07.
Scientific Objective: High accuracy land use/land cover mapping, forest type mapping and invasive shrub species location

Priority: a4h

PI: A. Zlinszky

Note: Data is currently symlinked from ~arsf/arsf_data/2010/flight_data/archive_overflow and when this is resolved the symlinks will need to be re-directed.

Sensors:

  • Eagle (19/05/2011)
  • Hawk (19/05/2011)
  • Leica LiDAR (04/03/2011)
  • Digital Photography (04/03/2011)

Change History (45)

comment:1 Changed 9 years ago by iopa

  • Description modified (diff)

comment:2 Changed 9 years ago by iopa

  • Description modified (diff)
  • Priority changed from alpha 5 to alpha 4 high

comment:3 Changed 9 years ago by iopa

The applanix and IPAS for this project have been copied from 238a. The processing need only be done once.

comment:4 Changed 9 years ago by iopa

Note; Eagle file VNIR238b-10-3.raw is still incomplete and needs to be recovered when ARSF Ops notify us that it is on thelma.

comment:5 Changed 9 years ago by knpa

  • Description modified (diff)

Edited scientific objective.

comment:6 Changed 9 years ago by knpa

  • Description modified (diff)

This flight will also be used for EUFAR10-07.

comment:7 Changed 9 years ago by iopa

From an email by Phil Goy on 19/10/2010,

Eagle file VNIR238b-10-3.raw is unrecoverable and we will have to process only the partial
one we already have.

The header for that still needs to be fixed to match the partial raw.

comment:8 Changed 9 years ago by iopa

The header has been fixed and the line is processable as any other, BUT

(for eagle line 3)
azspec_args_extra needs to contain (at least):
azspec_args-extra = -sdr eagle/darkFrames.bil -sdh eagle/darkFrames.bil.hdr

Those dark frames have been extracted from VNIR238b-10-1.raw, it has been tested as far as azspec and seems to work. If any problems arise in next stages for this line, keep this in mind.

comment:9 Changed 9 years ago by iopa

It turned out, as Phil and James told us when they visited, the above file was successfully recovered and then uploaded on thelma. After downloading and checking it, the file is complete and processable.

None of the above 'special care' will be necessary.

comment:10 Changed 9 years ago by knpa

Appears to be an issue with the bandset labels in the header files; they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:11 Changed 8 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Taking LIDAR.

comment:12 Changed 8 years ago by emca

Processed the camera imagery from raw to tiff

comment:13 Changed 8 years ago by knpa

Processed LIDAR. Images match-up well horizontally and vertically.

comment:14 Changed 8 years ago by knpa

Delivery created and ready for checking at: ~airborne/workspace/EUFAR10_01-2010_238b_Lake_Balaton_Peninsula/delivery/20110221/EUFAR10_01

comment:15 Changed 8 years ago by iopa

Checking LiDAR.

comment:16 Changed 8 years ago by iopa

LiDAR delivery check;

no errors found, data ready for delivery.

comment:17 Changed 8 years ago by emca

Processed camera data - ready for delivery check at
~airborne/workspace/EUFAR10_01-2010_238b_Lake_Balaton_Peninsula/delivery/20110302/EUFAR10_01

comment:18 Changed 8 years ago by anch

Starting camera delivery check

comment:19 Changed 8 years ago by anch

Camera delivery check complete.

All good and ready to deliver.

comment:20 Changed 8 years ago by anch

  • Description modified (diff)

Camera and LiDAR sent to PI

comment:21 Changed 8 years ago by anch

Starting hyperspectral processing.

comment:22 Changed 8 years ago by anch

  • Summary changed from EUFAR10/01, flight day 238b/2010, Lake Balaton Peninsula to EUFAR10/01 (AIMWETLAB), flight day 238b/2010, Lake Balaton Peninsula

comment:23 Changed 8 years ago by emca

LiDAR flightline 07 was missing from delivery (it was a copy of line 08). This line has been made available to the PI via FTP.

Updated DEM, screenshots, readme and missing line should be send with the next delivery to Zlinszky.

comment:24 Changed 8 years ago by iopa

Checking hyperspectral.

comment:25 Changed 8 years ago by iopa

  • Description modified (diff)

Hyperspectral checked and dispatched to Dr. Zlinszky on 19/05/11.

comment:26 Changed 8 years ago by knpa

SCTs (taken from .cfg file)

eagle hawk
1 0.03 0.00
2 0.03 0.00
3 0.89 1.57
4 0.00 0.00
5 0.00 0.00
6 0.00 0.03
7 0.04 0.00
8 0.03 -0.01
9 0.00 0.02
10 0.00 0.00
11 0.03 0.00
12 0.03 0.03
13 0.00 0.00

comment:27 Changed 8 years ago by knpa

Project r-synced, beginning archiving.

comment:28 Changed 8 years ago by knpa

AIMMS, GRIMM and Tabi data present.

comment:29 Changed 8 years ago by knpa

  • Component changed from Processing: general to Archiving

Tarball created, awaiting NEODC download (27/05/11)

comment:30 Changed 7 years ago by knpa

Backed-up by NEODC.

Moved to archive disk.

comment:31 Changed 7 years ago by knpa

Reprocessing hyperspectral with APL.

comment:32 Changed 7 years ago by knpa

Running stage 2.

comment:33 Changed 7 years ago by knpa

Re-running stage 2 (E+H) as apl chain has been updated to generate additional masked files.

comment:34 Changed 7 years ago by knpa

Re-running processing Hawk as aplcal has been edited again in relation to the bad pixel files.

In addition, Eagle 3, 4 and Hawk 4 didn't correctly process the first time.
Eagle raw 3 looks like it's broken, going to use VNIR238b-10-3_part.raw instead as that seems to have been used for the original delivery.
Line 4 appears to have no sync, not sure how it processed originally with an SCT value of 0. Am processing again and finding SCT manually.

comment:35 Changed 7 years ago by knpa

Running apl on VNIR238b-10-3_part.raw returns an error as the filesize is slightly off from header (about an 11th of a line).

Mistake in config when running for non-sync line 4 (e+h) so running again.

comment:36 Changed 7 years ago by knpa

Found SCTS for non-sync line 4:

e4: 1.37
h4: 0.43

comment:37 Changed 7 years ago by knpa

Removed non-reprocessing workspace version. Leaving ticket open for reprocessing.

comment:38 Changed 7 years ago by knpa

Fixed the above issue with eagle line 3 by truncating the line using head (creating VNIR238b-10-3_part_truncated.raw). This seems to have worked, now have 13 eagle lines that look good.

comment:39 Changed 7 years ago by knpa

Created delivery. Ready for delivery check.

comment:40 Changed 7 years ago by adbe

It looks like jaho delivery checked this project:

  • DEM should be named EUFAR10_01-2010_238b-ASTER.dem
  • The mapped eagle for line 3 only has 3 bands, should be ALL

comment:41 Changed 7 years ago by knpa

Corrected DEM name.

Re-generated e03 mapped file.

Ready to be delivered. (new mapped file needs to be zipped)

comment:42 Changed 7 years ago by anhi

Reprocessing hyperspectral delivered to Andras Zlinszky on 12/09/12 on disk 131.

comment:43 Changed 7 years ago by knpa

  • Milestone set to 2010 data processing completion

comment:44 Changed 7 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

comment:45 Changed 3 months ago by wja

Archiving
Created new lidar delivery with current structure.

Note: See TracTickets for help on using tickets.